# SSUMamba: Spatial-Spectral Selective State Space Model for Hyperspectral Image Denoising

### v1

@@ III-E Architecture Design Details
In the standard setting of our SSUMamba, we incorporate 6 VMamba blocks with SSAS with channel set as [32, 64, 64, 128, 128, 256]. The downsampling is enabled in block [2, 4, 6]. The SSUMamba is trained end-to-end to minimize the difference between the predicted clean HSI and the ground truth clean HSI. The loss function is set as:
@@
where denotes the HSI recovered from the noisy image in SSUMamba, denotes the corresponding clean image, and i the batch size within each iteration. An Adm optimizer with a learning rate of was used to train the SSUMamba model for 45 epochs and decreased with a factor of 0.3 in the 20th and 35th epoch. The batch size is set as 16. The model was implemented using PyTorch and trained on a single NVIDIA GeForce RTX 3090 GPU.
@@ IV Experiments
In this section, we present demonstrations highlighting the denoising prowess of the SSUMamba model quantitatively and qualitatively.
@@ IV-A1 Training and Testing Setting
In accordance with [11, 32], we opted for 100 HSIs from the Interdisciplinary Computational Vision Laboratory (ICVL) HSI dataset for training. These HSIs were captured using a Specim PS Kappa DX4 hyperspectral camera, sized at pixels and covering 31 spectral bands ranging from 400 to 700 nm. They were partitioned into patches measuring and subjected to augmentation through random flipping, cropping, and resizing. For testing, we assembled a the ICVL testing set and Houston 2018 HSI.
@@
| [0,15] | PSNR | 33.18 | 44.39 | 45.39 | 45.74 | 39.63 | 41.50 | 48.08 | 43.41 | 46.05 | 49.68 | 48.21 | 46.43 | 50.87 | 51.82 |
| SSIM | .6168 | .9683 | .9592 | .9657 | .8612 | .9356 | .9917 | .9315 | .9811 | .9912 | .9915 | .9878 | .9938 | .9948 | |
| SAM | .3368 | .0692 | .0845 | .0832 | .2144 | .1289 | .0404 | .0570 | .0560 | .0486 | .0387 | .0437 | .0298 | .0261 | |
| [0,55] | PSNR | 21.72 | 37.63 | 38.02 | 36.80 | 31.53 | 31.50 | 42.86 | 35.63 | 40.20 | 45.15 | 43.74 | 44.64 | 46.39 | 46.76 |
| SSIM | .2339 | .9008 | .8586 | .8285 | .6785 | .6233 | .9800 | .8125 | .9505 | .9810 | .9768 | .9840 | .9872 | .9875 | |
| SAM | .7012 | .1397 | .234 | .2316 | .4787 | .3583 | .063 | .1914 | .0993 | .0652 | .0582 | .0487 | .0457 | .0398 | |
| [0,95] | PSNR | 17.43 | 34.71 | 34.81 | 31.89 | 27.62 | 27.00 | 40.84 | 32.83 | 37.80 | 43.10 | 41.24 | 43.54 | 44.83 | 45.09 |
| SSIM | .1540 | .8402 | .7997 | .6885 | .5363 | .4208 | .9734 | .7482 | .9279 | .9734 | .9577 | .9806 | .9838 | .9837 | |
@@
| Mixture | PSNR | 13.21 | 23.36 | 27.55 | 18.23 | 23.61 | 23.10 | 27.58 | 30.93 | 34.90 | 34.09 | 28.44 | 39.73 | 39.22 | 41.63 |
| SSIM | .0841 | .4275 | .6743 | .1731 | .4448 | .3463 | .7250 | .8378 | .9041 | .9052 | .7393 | .9491 | .9626 | .9602 | |
| SAM | .9124 | .5476 | .5326 | .6873 | .6252 | .5144 | .4534 | .3613 | .1468 | .2340 | .4154 | .0869 | .0743 | .0798 | |
@@ IV-A2 Synthetic Data Generation
As indicated by [53], real-world noises in hyperspectral images (HSIs) typically follow non-independent and identical distribution (non-i.i.d.). Furthermore, remote-sensing HSIs commonly exhibit contamination from a blend of noises. Hence, in our experiments, we considered non-i.i.d. Gaussian noise and mixture noise. For non-i.i.d. Gaussian noise, we varied the standard deviation within the range of [0, 15], [0, 55], and [0, 95], and [0, 95]. The mixture noise comprised the following components: 1) non-i.i.d. Gaussian noise with [0,95]; 2) impulse noise affecting 1/3 of the bands, with intensities spanning from 10% to 70%; 3) strips affecting 5%-15% of columns on 1/3 of the bands; and 4) dead zones affecting 5%-15% of columns on 1/3 of the bands.
@@ IV-A4 Compared Methods
IV-B ICVL testing set
@@
The ICVL dataset comprises 50 HSIs from the broader ICVL dataset for synthetic testing. To alleviate the computational burden associated with certain model-based techniques like NGMeet, all these HSIs were resized to . Table I illustrates the quantitative analysis, with the top three results highlighted in bold red, bold blue, and bold green, respectively.
@@
FastHyDe demonstrates superior performance among model-based methods in Gaussian noise scenarios, attributed to its simultaneous modeling of the GSC and NSS of HSIs. For mixed noise scenarios, methods like LRTFL0 and E-3DTV exhibit improved performance. Deep-learning-based approaches leverage their data-driven learning capability, outperforming model-based methods consistently. T3SC and MAC-Net leverage physical modeling and data-driven learning, showcasing robustness against Gaussian noise. Constrained with spectral and spatial transformer-based modules, SST shows the best performance among all compared methods. Our SSUMamba generally outperforms SST. Even in mixed noise scenarios, where the transformer’s stronger modeling ability enables SST to outperform SSUMamba slightly in both SSIM and SAM metrics, we still outperform SST in PSNR by 2.17. In summary, SSUMamba stands out as the top performer in all assessed scenarios. This is attributed to its capability to adeptly capture the domain knowledge of HSIs through global spatial-spectral correlation modeling.
IV-C Houston 2018 HSI
| Model-based methods | Deep learning-based methods | ||||||||||||||
| [24] | [20] | [26] | [27] | [19] | [28] | [21] | [30] | [33] | [32] | [13] | [9] | (ours) | |||
| [0,15] | PSNR | 32.47 | 40.70 | 42.83 | 40.65 | 38.31 | 39.90 | 45.87 | 42.74 | 43.38 | 41.99 | 48.05 | 39.13 | 43.38 | 46.12 |
| SSIM | .6738 | .9667 | .9775 | .9468 | .8906 | .9601 | .9927 | .9842 | .9838 | .9850 | .9944 | .9779 | .9897 | .9929 | |
| SAM | .2562 | .0550 | .0357 | .0712 | .1279 | .0612 | .0355 | .0420 | .0429 | .0566 | .0270 | .0666 | .0401 | .0344 | |
| [0,55] | PSNR | 24.66 | 35.04 | 35.94 | 30.15 | 29.86 | 30.93 | 43.99 | 36.99 | 39.20 | 37.39 | 43.88 | 36.74 | 39.07 | 40.94 |
| SSIM | .3874 | .9002 | .8919 | .6390 | .6918 | .7624 | .9888 | .9386 | .9628 | .9670 | .9884 | .9714 | .9794 | .9830 | |
| SAM | .6484 | .1014 | .1387 | .3295 | .3707 | .1688 | .0391 | .1124 | .0698 | .0888 | .0393 | .0932 | .0710 | .0687 | |
| [0,95] | PSNR | 16.85 | 31.23 | 32.73 | 22.93 | 26.06 | 26.43 | 39.32 | 33.86 | 34.91 | 34.81 | 40.18 | 35.09 | 35.45 | 36.94 |
| SSIM | .2072 | .7766 | .8160 | .3724 | .5514 | .5662 | .9729 | .8730 | .9250 | .9466 | .9704 | .9559 | .9617 | .9630 | |
| SAM | .9201 | .1961 | .2160 | .5802 | .5400 | .3017 | .0715 | .1890 | .1116 | .1135 | .0542 | .1147 | .1178 | .0953 | |
| Mixture | PSNR | 11.72 | 22.76 | 25.86 | 15.58 | 22.36 | 21.84 | 27.07 | .1890 | .1116 | 29.84 | 28.83 | 32.55 | 31.07 | 34.23 |
| SSIM | .0843 | .4762 | .6933 | .1386 | .5169 | .3914 | .7757 | .8038 | .8570 | .8751 | .7963 | .9194 | .9166 | .9402 | |
| SAM | .9778 | .5168 | .4977 | .7652 | .5728 | .4857 | .4518 | .2221 | .1323 | .1943 | .2356 | .1241 | .1390 | .1103 | |
@@ IV-C Houston 2018 HSI
The Houston 2018 HSI was captured using the ITRES CASI hyperspectral imager, comprising pixels. The Houston HSI contains 48 bands spanning from 380 to 1050 nm. For our experiments, we used the final 46 bands of the Houston dataset, which is relatively clean, and extracted a cube from the center of the HSI for synthetic experiments. Deep learning models trained on the ICVL dataset were then employed for evaluation. However, it is noteworthy that T3SC, TRQ3D, and SST encounter challenges when dealing with HSIs that feature a distinct number of bands compared to the training data. To mitigate this issue, for T3SC, TRQ3D, and SST, we partitioned the provided HSI into multiple sub-images, each containing 31 bands for testing.
@@
The denoising outcomes are showcased in Table II. Given the substantial divergence between the Houston 2018 HSI and the training set ICVL, Deep-learning-based techniques face a formidable challenge when tasked with denoising and exhibit a decline in performace. Benefiting from the low-rank characteristic learned from data, MAC-Net achieve the best performance in Gaussian noise cases. However, it still suffuring from the limit adaptability to noise type, and get lower performance in mixture noise. Our proposed SSUMamba, despite experiencing some performance degradation in instances of significant data discrepancies, still outperforms transformer-based approaches consistently across all cases.
@@
Fig. 4 illustrates the reconstructed spatial images under mixed noise case. While conventional model-based approaches, tailored for Gaussian noise, struggle with eliminating artifacts such as stripes and artifacts, E-3DTV stands out by employing the -norm specifically for mixed noise scenarios. Due to spectral variations between the Houston dataset and the ICVL training set, TRQ3D and SST transformers exhibit some degree of failure in Houston image reconstruction. TRQ3D retaining Gaussian noise and stripe artifacts. SST exhibits color distortion, indicating inaccuracies in spectral fidelity relative to the original image, thus revealing limitations in accurately modeling spectral correlations. In contrast, our SSUMamba model utilizes alternating scan VMamba blocks to capture global spatial-spectral correlations effectively. This approach yields highly accurate reconstructions with minimal artifacts, particularly in scenarios with mixed noise, where detailed information tends to degrade. The subtle integration of long-range spatial-spectral information can help in the recovery of degraded data and substantially improve overall performance.

### v3

@@ III-D Architecture Design Details
In the standard setting of our SSUMamba, we incorporate 6 SSAS Mamba blocks with channel set as [32, 64, 64, 128, 128, 256]. The downsampling is enabled in block [2, 4, 6]. The SSUMamba is trained end-to-end to minimize the difference between the predicted clean HSI and the ground truth clean HSI. The loss function is set as:
@@
where denotes the HSI recovered from the noisy image in SSUMamba, denotes the corresponding clean image, and is the batch size within each iteration. An Adm optimizer with a learning rate of was used to train the SSUMamba model for 45 epochs and decreased with a factor of 0.5 in the 20th and 35th epochs. The batch size is set as 14. The model was implemented using PyTorch and trained on a single NVIDIA GeForce RTX 3090 GPU.
@@ IV Experiments
In this section, we demonstrate the denoising performance of the proposed SSUMamba through synthetic and real-world experiments.
@@ IV-A1 Training and Testing Setting
In accordance with [11, 33], we opted for 100 HSIs from the Interdisciplinary Computational Vision Laboratory (ICVL) HSI dataset for training. These HSIs were captured using a Specim PS Kappa DX4 hyperspectral camera, sized at pixels and covering 31 spectral bands ranging from 400 to 700 nm. They were partitioned into patches measuring and subjected to augmentation through random flipping, cropping, and resizing. The synthetic datasets include the ICVL testing set, Houston 2018 HSI, and Pavia city center HSI. The real-world datasets comprise the Gaofen-5 (GF-5) Wuhan HSI and the Earth Observing-1 (EO-1) HSI. The number of bands also increases from ICVL to EO-1 HSI, providing a comprehensive evaluation set to test all methods across varying numbers of bands.
@@
| [0,15] | PSNR | 33.18 | 44.39 | 45.39 | 45.74 | 39.63 | 41.50 | 48.08 | 43.41 | 46.05 | 49.68 | 48.21 | 46.43 | 50.87 | 51.47 |
| SSIM | .6168 | .9683 | .9592 | .9657 | .8612 | .9356 | .9917 | .9315 | .9811 | .9912 | .9915 | .9878 | .9938 | .9946 | |
| SAM | .3368 | .0692 | .0845 | .0832 | .2144 | .1289 | .0404 | .0570 | .0560 | .0486 | .0387 | .0437 | .0298 | .0252 | |
| [0,55] | PSNR | 21.72 | 37.63 | 38.02 | 36.80 | 31.53 | 31.50 | 42.86 | 35.63 | 40.20 | 45.15 | 43.74 | 44.64 | 46.39 | 46.82 |
| SSIM | .2339 | .9008 | .8586 | .8285 | .6785 | .6233 | .9800 | .8125 | .9505 | .9810 | .9768 | .9840 | .9872 | .9881 | |
| SAM | .7012 | .1397 | .234 | .2316 | .4787 | .3583 | .0630 | .1914 | .0993 | .0652 | .0582 | .0487 | .0457 | .0371 | |
| [0,95] | PSNR | 17.43 | 34.71 | 34.81 | 31.89 | 27.62 | 27.00 | 40.84 | 32.83 | 37.80 | 43.10 | 41.24 | 43.54 | 44.83 | 45.50 |
| SSIM | .1540 | .8402 | .7997 | .6885 | .5363 | .4208 | .9734 | .7482 | .9279 | .9734 | .9577 | .9806 | .9838 | .9852 | |
@@
| Mixture | PSNR | 13.21 | 23.36 | 27.55 | 18.23 | 23.61 | 23.10 | 27.58 | 30.93 | 34.90 | 34.09 | 28.44 | 39.73 | 39.22 | 42.76 |
| SSIM | .0841 | .4275 | .6743 | .1731 | .4448 | .3463 | .7250 | .8378 | .9041 | .9052 | .7393 | .9491 | .9626 | .9701 | |
| SAM | .9124 | .5476 | .5326 | .6873 | .6252 | .5144 | .4534 | .3613 | .1468 | .2340 | .4154 | .0869 | .0743 | .0710 | |
@@ IV-A2 Synthetic Data Generation
As indicated by [56], real-world noises in hyperspectral images (HSIs) typically follow non-independent and identical distribution (non-i.i.d.). Furthermore, remote-sensing HSIs commonly exhibit contamination from a mixture of noises. Hence, in our experiments, we considered non-i.i.d. Gaussian noise and mixture noise. For non-i.i.d. Gaussian noise, we varied the standard deviation within the range of , and . The mixture noise comprised the following components: 1) non-i.i.d. Gaussian noise with ; 2) impulse noise affecting 1/3 of the bands, with intensities spanning from 10% to 70%; 3) strips affecting 5%-15% of columns on 1/3 of the bands; and 4) dead zones affecting 5%-15% of columns on 1/3 of the bands.
@@ IV-A4 Compared Methods
IV-B Comparison on Synthetic Datasets
To demonstrate the denoising performance both quantitatively and qualitatively, we first present the denoising results on the ICVL testing set, Houston 2018, and Pavia city center HSIs.
IV-B1 ICVL testing set
@@
The ICVL dataset comprises 50 HSIs from the ICVL dataset for synthetic testing. To alleviate the computational burden associated with certain model-based techniques like NGMeet, all these HSIs were resized to . The ICVL dataset is enriched with a greater number of HSIs, offering a comprehensive evaluation platform to assess the performance of all methods under different noise cases. Thus, both Gaussian and mixture noises were included for assessment. Table I illustrates the quantitative analysis, with the top three results highlighted in bold red, bold blue, and bold green, respectively.
@@
FastHyDe demonstrates superior performance among model-based methods in Gaussian noise cases, attributed to its simultaneous modeling of the GSC and NSS of HSIs. For mixture noise cases, methods like LRTF and E-3DTV exhibit improved performance. Deep-learning-based approaches leverage their data-driven learning capability, outperforming model-based methods consistently. T3SC and MAC-Net leverage physical modeling and data-driven learning, showcasing robustness against Gaussian noise. Constrained with spectral and spatial transformer-based modules, SST shows the best performance among all compared methods. In summary, SSUMamba stands out as the top performer in all assessed cases. This is attributed to its capability to adeptly capture the domain knowledge of HSIs through global spatial-spectral correlation modeling.
IV-B2 Houston 2018 HSI
| Model-based methods | Deep learning-based methods | |||||||||||||
| Index | Noisy | BM4D | MTSNMF | LLRT | NGMeet | LRMR | FastHyDe | LRTF | E-3DTV | T3SC | MAC-Net | TRQ3D | SST | SSUMamba |
| [25] | [21] | [27] | [28] | [20] | [29] | [22] | [31] | [34] | [33] | [13] | [9] | (ours) | ||
| PSNR | 11.72 | 22.76 | 25.86 | 15.58 | 22.36 | 21.84 | 27.07 | 28.75 | 30.64 | 29.84 | 28.83 | 32.55 | 31.07 | 34.55 |
| SSIM | .0843 | .4762 | .6933 | .1386 | .5169 | .3914 | .7757 | .8038 | .8570 | .8751 | .7963 | .9194 | .9166 | .9444 |
| SAM | .9778 | .5168 | .4977 | .7652 | .5728 | .4857 | .4518 | .2221 | .1323 | .1943 | .2356 | .1241 | .1390 | .1047 |
@@ IV-B2 Houston 2018 HSI
The Houston 2018 HSI was captured using the ITRES CASI hyperspectral imager, comprising pixels. The Houston HSI contains 48 bands spanning from 380 to 1050 nm. For our experiments, we omit the first 2 bands, which is relatively noisy, and extracted a cube from the center of the HSI for synthetic experiments. To replicate real-world noise found in remote-sensing images, we introduced mixture noise to the pristine HSI, resulting in a noisy HSI. Deep learning models trained on the ICVL dataset were then employed for evaluation. However, it is noteworthy that T3SC, TRQ3D, and SST encounter challenges when dealing with HSIs that feature a distinct number of bands compared to the training data. To mitigate this issue, for T3SC, TRQ3D, and SST, we partitioned the provided HSI into multiple sub-images, each containing 31 bands for testing.
| Model-based methods | Deep learning-based methods | |||||||||||||
| Index | Noisy | BM4D | MTSNMF | LLRT | NGMeet | LRMR | FastHyDe | LRTF | E-3DTV | T3SC | MAC-Net | TRQ3D | SST | SSUMamba |
| [25] | [21] | [27] | [28] | [20] | [29] | [22] | [31] | [34] | [33] | [13] | [9] | (ours) | ||
| PSNR | 13.46 | 21.70 | 25.15 | 15.44 | 23.68 | 22.72 | 26.78 | 26.49 | 30.44 | 28.69 | 27.74 | 28.23 | 31.87 | 35.58 |
| SSIM | .2181 | .5128 | .7397 | .2760 | .7176 | .6524 | .8361 | .8147 | .8941 | .8656 | .8724 | .8665 | .9234 | .9557 |
| SAM | .8893 | .5297 | .4094 | .7902 | .4760 | .3771 | .4040 | .3703 | .1134 | .2135 | .3222 | .1961 | .1676 | .1055 |
@@
The denoising results are showcased in Table II. All deep learning-based methods learn the mixture noise cases in the ICVL dataset and perform better than model-based methods. However, given the substantial divergence between the Houston 2018 HSI and the training set ICVL, deep learning-based techniques face a formidable challenge when tasked with denoising and exhibiting a decline in performance. Our proposed SSUMamba, despite experiencing some performance degradation in instances of significant data discrepancies, still outperforms transformer-based approaches.
@@
Fig. 4 illustrates the false-color images recovered from all compared methods. While conventional model-based approaches, tailored for Gaussian noise, struggle with eliminating artifacts such as stripes and artifacts, E-3DTV stands out by employing the -norm for noise in the mixture noise. Due to spectral variations between the Houston HSI and the ICVL training set, TRQ3D and SST transformers exhibit some degree of failure in Houston image reconstruction. TRQ3D retaining Gaussian noise and stripe artifacts. SST exhibits color distortion, indicating inaccuracies in spectral fidelity relative to the original image, thus revealing limitations in accurately modeling spectral correlations. In contrast, our SSUMamba model utilizes alternating scan Mamba blocks to capture global spatial-spectral correlations effectively. Our method yields highly accurate reconstructions with minimal artifacts, particularly in cases with mixture noise, where detailed information degrades. Exploiting long-range spatial-spectral information can help recover degraded data and substantially improve overall performance.
The spectral reflectances are shown in Fig. 5, where the mixture noise makes the spectral distortion drastic, as in Fig. 5(b). Because model-based methods are susceptible to noise type, they struggle to generate recovered spectra that match clean image spectra in denoising mixture noise. Taking into account the mixed noise, E-3DTV models the low-rankness of the spatial and spectral gradient map in HSI, resulting in the best match with the clean HSI among model-based methods. Deep learning-based T3SC and E-3DTV design for Gaussian noise recover the spectral reflectance with jitters. TRQ3D and SST cannot accommodate the processing as the number of bands in Houston 2018 differs from that in the ICVL training set, resulting in incomplete GSC and less accurate matching of spectral reflectances with clean HSI. Our SSUMamba, with the ability to capture global spatial-spectral correlations, can effectively recover spectral reflectance, demonstrating the best performance in spectral modeling ability among all methods.
IV-B3 Pavia City Center HSI
The Pavia City Center HSI was captured using the Reflective Optics System Imaging Spectrometer (ROSIS) sensor spanning from 430 to 860 nm. Following [57], we removed the first 22 noisy bands and added mixture noise on a subimage with 80 bands for synthetic experiments. Other experimental settings were maintained in line with those used for the Houston 2018 HSI.
Table III shows a quantitative presentation of the results of the experiment. Model-based E-3DTV calculates sparsity on gradient map subspaces across all HSI bands. Its ability to better reflect spectral correlations and differences allows E-3DTV to outperform TRQ3D, which introduces RNNs to model incomplete spectral correlations in the HSI, especially when the Pavia City Center HSI has different spectral features from the training data. Exploiting long-range correlations in spectral and spatial dimensions, transformer-based SST shows better performance than TRQ3D. However, it still struggles when dealing with testing data with a different number of bands. Our SSUMamba achieves the best performance in the Pavia City Center HSI by exploiting complete global spatial-spectral correlations with SSM.
Fig. 6 shows the false-color images recovered by all compared methods. The model-based methods, designed for Gaussian noise, fail to eliminate artifacts such as stripes and artifacts in the mixture noise. E-3DTV, which models the low-rankness of the spatial and spectral gradient map in HSI, achieves the best performance in the mixture noise. However, it still gets blurred due to the limited modeling ability in spatial. Transformer-based methods, such as TRQ3D and SST, exhibit some degree of failure in Pavia City Center HSI reconstruction due to spectral variations between the Pavia City Center HSI and the training set ICVL. Our SSUMamba, which captures global spatial-spectral correlations effectively, yields highly accurate reconstructions with minimal artifacts, particularly in cases with heavy noise, where detailed information degrades. Fig. 7 shows the spectral reflectances of the mixture noise. The proposed SSUMamba effectively exploits global spatial-spectral correlations to recover the spectral reflectance, resulting in the best match with the clean HSI among all methods.
IV-C Comparison on Real-World HSIs
We carried out extra experiments on real-world remote-sensing HSIs to thoroughly evaluate the effectiveness of denoising across various methods. We provide a qualitative comparative analysis to assess the denoising results.
IV-C1 Gaofen-5 Wuhan HSI
The Gaofen-5 (GF-5) Wuhan HSI was acquired using the Advanced Hyperspectral Imager (AHSI) with a resolution of pixels and 155 spectral bands. In our real-world experiment, we applied all compared methods to the original noisy HSI. The results are presented in Fig. 8, where the original image is degraded by impulse and stripes. Among model-based methods, FastHyDe and LRTF achieved better performance thanks to their ability to recover from mixture noise. T3SC and MAC-Net effectively removed most of the stripes in the GF-5 Wuhan HSI. T3SC produced sharper results by learning both spectral and spatial information in a trainable sparse coding model. However, Transformer-based TRQ3D and SST failed to completely remove all the stripes due to limited adaptability to the imaging environment. Our SSUMamba method, which effectively captures global spatial-spectral correlations to learn the deep 3-D features of the HSI, yielded highly accurate reconstructions with minimal artifacts and sharper edges compared to all other methods.
Spectral reflectances are depicted in Fig. 7. The heavy impulse and stripes cause significant spectral distortion, as seen in Fig. 9(a). T3SC and SST process the HSI in patches, leading to incomplete global spatial-spectral correlations and spectral jitters. In contrast, our SSUMamba, which captures global spatial-spectral correlations effectively, can recover spectral reflectance more accurately, resulting in smoother curves that closely resemble the original shape.
IV-C2 Earth Observing-1 HSI
The Earth Observing-1 (EO-1) HSI was acquired by Hyperion with 242 bands and a resolution of pixels. Following [20], the water absorption bands were removed, and the final HSI was cropped to pixels with 166 bands. Similar to the experiment with the GF-5 Wuhan HSI, we applied all compared methods to the original noisy HSI. The results are presented in Fig. 10. The presence of heavier impulse and stripe noise makes the denoising task more challenging. Besides E-3DTV, all other model-based methods failed to effectively remove the stripes and impulse noise due to the heavier noise. Although E-3DTV removed the stripes and impulse noise, it resulted in blurred edges. Among deep learning-based methods, T3SC and TRQ3D successfully removed most of the stripes. However, insufficient modeling of global spatial-spectral correlations resulted in blurring similar to E-3DTV. MAC-Net and SST were unable to remove all the stripes, while SST introduced color distortion. Our SSUMamba effectively learned the deep 3-D nature of the HSI, incorporating spatial and spectral information to produce sharper edges and more accurate colors. Spectral reflectances are shown in Fig. 11, where jitters and discontinuities are evident in Fig. 11(a). SSUMamba achieved smoother and more continuous spectral reflectance, closer to the clean HSI.